# Retrieval-Augmented Decision Transformer: External Memory for In-context RL

### v2

@@ Abstract
In-context learning (ICL) is the ability of a model to learn a new task by observing a few exemplars in its context. While prevalent in NLP, this capability has recently also been observed in Reinforcement Learning (RL) settings. Prior in-context RL methods, however, require entire episodes in the agent’s context. Given that complex environments typically lead to long episodes with sparse rewards, these methods are constrained to simple environments with short episodes. To address these challenges, we introduce Retrieval-Augmented Decision Transformer (RA-DT). RA-DT employs an external memory mechanism to store past experiences from which it retrieves only sub-trajectories relevant for the current situation. The retrieval component in RA-DT does not require training and can be entirely domain-agnostic. We evaluate the capabilities of RA-DT on grid-world environments, robotics simulations, and procedurally-generated video games. On grid-worlds, RA-DT outperforms baselines, while using only a fraction of their context length. Furthermore, we illuminate the limitations of current in-context RL methods on complex environments and discuss future directions. To facilitate future research, we release datasets for four of the considered environments.
@@ 1 Introduction
In-context Learning (ICL) is the ability of a model to learn new tasks by leveraging a few exemplars in its context [Brown et al., 2020]. Large Language Models (LLMs) exhibit this capability after pre-training on large amounts of data crawled from the web. A similar trend has emerged in the field of RL, where agents are pre-trained on datasets with an increasing number of tasks [Chen et al., 2021; Janner et al., 2021; Reed et al., 2022; Lee et al., 2022; Brohan et al., 2022; 2023]. After training, such an agent is capable of learning new tasks by observing previous trials in its context [Laskin et al., 2022; Liu & Abbeel, 2023; Lee et al., 2023; Raparthy et al., 2023]. Consequently, ICL is a promising direction for generalist agents to acquire new tasks without the need for re-training, fine-tuning, or providing expert-demonstrations.
@@
Existing methods for in-context RL rely on keeping entire episodes in their context [Laskin et al., 2022; Lee et al., 2023; Kirsch et al., 2023; Raparthy et al., 2023]. Consequently, these methods face challenges in complex environments, as complex environments are usually characterized by long episodes and sparse rewards. Episodes in RL may consist of thousands of interaction steps, and processing them is computationally expensive, especially for network architectures such as the Transformer [Vaswani et al., 2017]. Furthermore, not all information an agent encountered in the past may be necessary to solve the new task. Therefore, we address the question of how to facilitate ICL for environments with long episodes and sparse rewards.
@@
Figure 1: Illustration of Retrieval-augmented Decision Transformer (RA-DT). Left: Prior to training, we encode pre-collected trajectories via an embedding model. During training, we retrieve sub-trajectories using the current context as query, and fuse them into layers via cross-attention. Right: During inference, the collected experience is stored and retrieved during environment interaction.
@@
We introduce Retrieval-Augmented Decision Transformer (RA-DT), which incorporates an external memory into the Decision Transformer [Chen et al., 2021, DT] architecture (see Figure 1). Our external memory enables efficient storage and retrieval of past experiences, that are relevant for the current situation. We achieve this by leveraging a vector index populated with sub-trajectories, in combination with maximum inner product search; akin to Retrieval-augmented Generation (RAG) in LLMs [Khandelwal et al., 2019; Lewis et al., 2020; Borgeaud et al., 2022]. To encode retrieved sub-trajectories, RA-DT relies on a pre-trained embedding model, which can either be domain-specific, such as a DT trained on the same domain, or a domain-agnostic language model (LM) (see Section 3). Subsequently, RA-DT uses cross-attention to leverage the retrieved sub-trajectories and predict the next action. This way, RA-DT does not rely on a long context and can deal with sparse reward settings.
@@
We evaluate the effectiveness of RA-DT on grid-world environments used in prior work with sparse rewards and increasing grid-sizes (Dark-Room, Dark Key-Door, Maze-Runner), robotics environments (Meta-World, DMControl) and procedurally-generated video games (Procgen). On grid-worlds, RA-DT considerably outperforms previous in-context RL methods, while only using a fraction of their context length. Further, we show that our domain-agnostic trajectory embedding model reaches performance close to a domain-specific one. On the remaining more complex environments, we observe consistent improvements for RA-DT on hold-out tasks, but no in-context improvement for any method. Therefore, we discuss the current limitations of RA-DT and other in-context RL methods and elaborate on potential remedies and future directions for in-context RL.
@@
  • We introduce Retrieval-augmented Decision Transformers (RA-DT) and evaluate its effectiveness on a number of diverse domains.
  • We show that a domain-agnostic embedding model can be utilized for retrieval in RL without requiring any pre-training, and reaches performance close to a domain-specific model.
  • We release datasets for Dark-Room, Dark Key-Door, Maze-Runner, and Procgen to foster future research on in-context decision-making that leverages offline pre-training. [^1] GitHub: https://github.com/ml-jku/RA-DT.
@@ 2 Related Work
In-context Learning. ICL is a form of Meta-learning, also referred to as learning-to-learn [Schmidhuber, 1987]. Typically, meta-learning is targeted and learned through a meta-training phase, for example in supervised-learning [Santoro et al., 2016; Mishra et al., 2018; Finn et al., 2017] or in RL [Wang et al., 2016; Duan et al., 2016; Kirsch et al., 2019; Flennerhag et al., 2019]. In contrast, ICL emerges as a result of pre-training on a certain data distribution [Chan et al., 2022]. This ability was first observed in Hochreiter et al. [2001] via LSTMs [Hochreiter & Schmidhuber, 1997] and later re-discovered in LLMs [Brown et al., 2020]. Ortega et al. [2019] found that every memory-based architecture may exhibit such capabilities. Another crucial factor is a training distribution comprising a vast amount of tasks [Chan et al., 2022; Kirsch et al., 2022]. Recent works combined these properties to induce ICL in RL [Laskin et al., 2022; Lee et al., 2022; Kirsch et al., 2023]. While promising, they require keeping entire episodes in context, which is difficult in environments with long episodes. Raparthy et al. [2023] consider an in-context imitation learning setting given expert demonstrations. In contrast, RA-DT can handle long episodes and does not rely on expert demonstrations.
@@
Retrieval-augmented Generation. The aim of retrieval-augmentation is to provide a model access to an external memory. This alleviates the need to store the training data in the parameters of a model and allows to condition on new data without re-training. RAG is successfully applied in the realm of LLMs [Khandelwal et al., 2019; Guu et al., 2020; Lewis et al., 2020; Borgeaud et al., 2022; Izacard et al., 2022; Ram et al., 2023], multi-modal language generation [Hu et al., 2023; Yasunaga et al., 2023; Yang et al., 2023b; Ramos et al., 2022], and for chemical reaction prediction [Seidl et al., 2022]. In RL, the access to an external memory is often referred to as episodic memory [Sprechmann et al., 2018; Blundell et al., 2016; Pritzel et al., 2017]. Goyal et al. [2022] investigate the effect of different data sources in the external memory of an online RL agent. [Humphreys et al., 2022] provide access to millions of expert demonstrations via RAG in the game of Go. In contrast, RA-DT does not rely on expert demonstrations, but leverages RAG to learn new tasks entirely in-context without the need for weight updates. Further, RA-DT does not rely on a pre-trained domain-specific embedding model, as we demonstrate that the embedding model can be entirely domain-agnostic.
@@
External memory in RL. Most prior works have explored the utility of an external memory to cope with partially observable environments [Åström, 1965; Kaelbling et al., 1998], in which the agent must remember past events to approximate the true state of the environment. This is difficult, especially for complex tasks with sparse rewards [Arjona-Medina et al., 2019; Patil et al., 2022; Widrich et al., 2021] and long episodes. To cope with this problem, Neural Turing Machines [Graves et al., 2014], which rely on a neural controller to read from and write to an external memory, were applied to RL [Zaremba & Sutskever, 2015]. Memory networks [Weston et al., 2015] leverage an external memory for reasoning. Wayne et al. [2018] propose a memory architecture with read/write access to learn what information to store based on a world model. In contrast, RA-DT only retrieves pieces of past information similar to the current encountered situation. Hill et al. [2021] propose an attention-based external memory, where queries, keys, and values are represented by different modalities. Similarly, our domain-agnostic embedding model extends the idea of history compression via LLMs [Paischer et al., 2022; 2023] to retrieval, where queries and keys are encoded in the language space, while values comprise raw sub-trajectories.
@@ 3.1 Background
Decision Transformer. Decision Transformer [Chen et al., 2021, DT] learns a policy from offline data by conditioning on future rewards. This allows rephrasing RL as a sequence modelling problem, where the agent is trained in a supervised manner to map future rewards to actions, often referred to as upside-down RL [Schmidhuber, 2019]. To train the DT, we assume access to a pre-collected dataset $\mathcal{D}=\{\tau_{i}\mid 1\leq i\leq N\}$ of $N$ trajectories $\tau_{i}$ that are sampled from the environment via a behavioural policy $\pi_{\beta}$. Each trajectory $\tau\in\mathcal{D}$ consists of state, action, reward, and return-to-go (RTG) quadruplets $\tau_{i}=(s_{0},a_{0},r_{0},\hat{R}_{0},\ldots,s_{T},a_{T},r_{T},\hat{R}_{T})$, where $T$ represents the length of trajectory $\tau_{i}$, and $\hat{R}_{t}=\sum_{t^{\prime}=t}^{T}r_{t^{\prime}}$. The DT $\pi_{\theta}$ is trained to predict the ground truth action $a_{t}$ conditioned on sub-trajectories via cross-entropy or mean-squared error loss, depending on the domain:
@@ 3.2 Retrieval-augmented Decision Transformer (RA-DT)
Figure 2: Illustration of experience reweighting. Given a query trajectory, we retrieve the top $l>k$ most relevant experiences by maximum inner product search. Each experience has an associated task ID, and return, based on which we compute their utility. We reweight by $s_{rel}$ and $s_{u}$, to obtain the final retrieval score $s_{ret}$, and return the top-$k$ experiences.
Processing long sequences with DTs is computationally expensive due to the quadratic complexity of the Transformer architecture. To address this challenge, we introduce RA-DT, which equips the DT with an external memory that relies on a vector index for retrieval. Consequently, RA-DT consists of a parametric and a non-parametric component, reminiscent of complementary learning systems [Mcclelland et al., 1995; Kumaran et al., 2016]. The former is represented by the DT and learns to predict actions conditioned on the future return. The latter is the retrieval component that searches for relevant experiences, similar to Borgeaud et al. [2022] (see Figure 1).
@@ 3.2.1 Vector Index for Retrieval Augmentation
We aim at augmenting the DT with a vector index (external memory) that allows for retrieval of relevant experiences. To this end, we build our vector index by leveraging an embedding model $g:\tau\mapsto\mathbb{R}^{d_{r}}$ that takes a trajectory $\tau$ and returns a vector of size $d_{r}$. Given a dataset $\mathcal{D}$ of trajectories, we obtain a set of key-value pairs of our vector index by embedding all sub-trajectories $\tau_{t-C:t}\in\mathcal{D}$ via $g(\cdot)$ to obtain $\mathcal{K}\times\mathcal{V}=\{(g(\tau_{i,t-C:t}),\tau_{i,t-C:t+C})\mid 1\leq i\leq|\mathcal{D}|\}$. Note that values contain sub-trajectories ranging from $t-C$ to $t+C$, while keys use sub-trajectories $t-C:t$ for a fixed $C$, where $t$ goes over trajectory length in increments of $C$ (see Appendix C.4 for more details). The reason for this choice is that during inference, the model does not have access to future states.
In RAG applications for Natural Language Processing (NLP), a common choice for $g(\cdot)$ is a pre-trained LM. While pre-trained models in NLP are ubiquitous, they are rarely available in RL. A natural choice to instantiate $g(\cdot)$ is to train a DT on the pre-collected dataset $\mathcal{D}$, as they exhibit a well-separated embedding space after pre-training [Schmied et al., 2024]. Therefore, they are well suited for retrieval since a new task can be matched to similar tasks in the vector index. As a domain-agnostic alternative, we propose to utilize the FrozenHopfield (FH) mechanism Paischer et al. [2022] to map trajectories to the embedding space of a pre-trained LM. This enables instantiating $g(\cdot)$ with a pre-trained language encoder. The FH mechanism is parameterized by an embedding matrix $\bm{E}\in\mathbb{R}^{v\times d_{\text{LM}}}$ of a pretrained LM with vocabulary size $v$ and hidden dimension $d_{\text{LM}}$, a random matrix $\bm{P}$ with entries sampled from $\mathcal{N}(0,d_{\text{in}}/d_{\text{LM}})$, and a scaling factor $\beta$ and performs:
@@
We denote $\bm{x}_{t}\in\mathbb{R}^{d_{\text{in}}}$ as the input token and apply the FH position-wise to every state/action/reward token in a sub-trajectory $\tau_{t-C:t}$ separately. Finally, we apply a LM on top of the FH to obtain the keys of our vector index by setting $g(\cdot)=\operatorname{LM}(\operatorname{FH}(\cdot))$. Utilizing the FH enables leveraging the expressive power of pre-trained LMs as trajectory encoders for RL. This sidesteps the need for pre-training a domain-specific model and can be incorporated in any existing retrieval-augmentation pipeline.

### v3

@@ Abstract
In-context learning (ICL) is the ability of a model to learn a new task by observing a few exemplars within its context. While prevalent in NLP, this capability has recently also been observed in Reinforcement Learning (RL) settings. Prior in-context RL methods, however, require entire episodes in the agent’s context. Given that complex environments typically lead to long episodes with sparse rewards, these methods are constrained to environments with short episodes. To address these challenges, we introduce Retrieval-Augmented Decision Transformer (RA-DT). RA-DT employs an external memory mechanism to store past experiences from which it retrieves only sub-trajectories relevant for the current situation. The retrieval component in RA-DT can be entirely domain-agnostic. We evaluate the capabilities of RA-DT on grid-world environments, robotics simulations, and procedurally-generated video games. On grid-worlds, RA-DT outperforms baselines while using only a fraction of their context length. Furthermore, we illuminate the limitations of current in-context RL methods on complex environments and discuss future directions. To facilitate future research, we release datasets for four of the considered environments [^1] GitHub: https://github.com/ml-jku/RA-DT.
@@ 1 Introduction
In-context Learning (ICL) is the ability of a model to learn new tasks by leveraging a few exemplars within its context (Brown et al., 2020). Large Language Models (LLMs) exhibit this capability after pre-training on large amounts of data crawled from the web. A similar trend has emerged in the field of RL, where agents are pre-trained on datasets with an increasing number of tasks (Chen et al., 2021; Janner et al., 2021; Reed et al., 2022; Lee et al., 2022; Brohan et al., 2022; 2023). After training, such an agent is capable of learning new tasks by observing previous trials in its context (Laskin et al., 2022; Liu & Abbeel, 2023; Lee et al., 2023; Raparthy et al., 2023). Consequently, ICL is a promising direction for generalist agents to acquire new tasks without the need for re-training, fine-tuning, or providing expert demonstrations.
@@
Existing methods for in-context RL rely on keeping entire episodes in their context (Laskin et al., 2022; Lee et al., 2023; Kirsch et al., 2023; Raparthy et al., 2023). Consequently, these methods face challenges in environments with long episodes and sparse rewards, which are ubiquitous in RL. Episodes in RL may consist of thousands of interaction steps (e.g., Atari or real-world scenarios), and processing them is computationally expensive, especially for network architectures such as the Transformer (Vaswani et al., 2017). Furthermore, not all information an agent encountered in the past may be necessary to solve the new task. Therefore, we address the question of how to facilitate ICL for environments with long episodes and sparse rewards.
@@
Figure 1: Illustration of Retrieval-augmented Decision Transformer (RA-DT). Left: Prior to training, we encode pre-collected trajectories via an embedding model. During training, we retrieve sub-trajectories using the current context as a query and fuse them into layers via cross-attention. Right: During inference, the collected experiences are stored in a retrieval buffer and selectively retrieved during environment interaction depending on the current situation.
@@
We introduce Retrieval-Augmented Decision Transformer (RA-DT), which incorporates an external memory into the Decision Transformer (Chen et al., 2021, DT) architecture (see Figure 1). Our external memory enables efficient storage and retrieval of past experiences that are relevant to the current situation. We achieve this by leveraging a vector index populated with sub-trajectories, in combination with maximum inner product search; akin to Retrieval-augmented Generation (RAG) in LLMs (Khandelwal et al., 2019; Lewis et al., 2020; Borgeaud et al., 2022). To encode retrieved sub-trajectories, RA-DT relies on a pre-trained embedding model, which can either be domain-specific, such as a DT trained on the same domain, or a domain-agnostic off-the-shelf language model (LM) (see Section 3). Subsequently, RA-DT uses cross-attention to leverage the retrieved sub-trajectories and predicts the next action. This way, RA-DT does not rely on an excessively long context and can handle sparse reward settings.
@@
We evaluate the effectiveness of RA-DT on grid-world environments used in prior work with sparse rewards and increasing grid-sizes (Dark-Room, Dark Key-Door, Maze-Runner), robotics environments (Meta-World, DMControl), and procedurally-generated video games (Procgen). On grid-worlds, RA-DT considerably outperforms previous in-context RL methods, while only using a fraction of their context length. Furthermore, we demonstrate that our domain-agnostic trajectory embedding model achieves performance similar to a domain-specific one. On the remaining more complex environments (robotics and video games), we observe consistent performance gains for RA-DT on hold-out tasks, but no in-context improvement for any method. Therefore, we discuss the current limitations of RA-DT and other in-context RL methods and elaborate on potential remedies and future directions for in-context RL.
@@
  • We introduce Retrieval-augmented Decision Transformer (RA-DT) and evaluate its effectiveness on a variety of diverse domains, including grid-worlds, video games, and robotics environments.
  • We demonstrate that our domain-agnostic embedding model can be utilized for retrieval in RL without requiring pre-training on the target domain, and achieves performance close to a domain-specific model.
  • We release datasets for Dark-Room, Dark Key-Door, Maze-Runner, and Procgen to foster future research on in-context decision-making that leverages offline pre-training.
@@ 2 Related Work
In-context Learning. ICL is a form of Meta-learning, also referred to as learning-to-learn (Schmidhuber, 1987). Typically, meta-learning is targeted and learned through a meta-training phase, for example, in supervised learning (Santoro et al., 2016; Mishra et al., 2018; Finn et al., 2017) or in RL (Wang et al., 2016; Duan et al., 2016; Kirsch et al., 2019; Flennerhag et al., 2019). In contrast, ICL emerges as a result of pre-training on a certain data distribution (Chan et al., 2022). This ability was first observed in Hochreiter et al. (2001) via LSTMs (Hochreiter & Schmidhuber, 1997) and later re-discovered in LLMs (Brown et al., 2020). Ortega et al. (2019) found that every memory-based architecture may exhibit such capabilities. Another crucial factor is a training distribution comprising a vast amount of tasks (Chan et al., 2022; Kirsch et al., 2022). Recent works combined these properties to induce ICL in RL (Laskin et al., 2022; Lee et al., 2022; Kirsch et al., 2023). While promising, they require keeping entire episodes in context, which is difficult in environments with long episodes. Raparthy et al. (2023) consider an in-context imitation learning setting given expert demonstrations. In contrast, RA-DT can handle long episodes and does not rely on expert demonstrations.
@@
Retrieval-augmented Generation. The aim of retrieval-augmentation is to provide a model access to an external memory. This alleviates the need to store the training data in the parameters of a model and allows conditioning on new data without re-training. RAG is successfully applied in the realm of LLMs (Khandelwal et al., 2019; Guu et al., 2020; Lewis et al., 2020; Borgeaud et al., 2022; Izacard et al., 2022; Ram et al., 2023), multi-modal language generation (Hu et al., 2023; Yasunaga et al., 2023; Yang et al., 2023; Ramos et al., 2022), and for chemical reaction prediction (Seidl et al., 2022). In RL, the access to an external memory is often referred to as episodic memory (Sprechmann et al., 2018; Blundell et al., 2016; Pritzel et al., 2017). Goyal et al. (2022) investigate the effect of different data sources in the external memory of an online RL agent. (Humphreys et al., 2022) provide access to millions of expert demonstrations via RAG in the game of Go. In contrast, RA-DT does not rely on expert demonstrations but leverages RAG to learn new tasks entirely in context without the need for weight updates. Furthermore, RA-DT does not rely on a pre-trained domain-specific embedding model, as we demonstrate that the embedding model can be entirely domain-agnostic.
@@
External memory in RL. Most prior works have explored the utility of an external memory to cope with partially observable environments (Åström, 1965; Kaelbling et al., 1998), in which the agent must remember past events to approximate the true state of the environment. This is difficult, especially for complex tasks with sparse rewards (Arjona-Medina et al., 2019; Patil et al., 2022; Widrich et al., 2021) and long episodes. To cope with this problem, Neural Turing Machines (Graves et al., 2014), which rely on a neural controller to read from and write to an external memory, were applied to RL (Zaremba & Sutskever, 2015). Memory networks (Weston et al., 2015) leverage an external memory for reasoning. Wayne et al. (2018) propose a memory architecture with read/write access to learn what information to store based on a world model. In contrast, RA-DT only retrieves pieces of past information similar to the current encountered situation. Hill et al. (2021) propose an attention-based external memory, where queries, keys, and values are represented by different modalities. Similarly, our domain-agnostic embedding model extends the idea of history compression via LLMs (Paischer et al., 2022; 2023) to retrieval, where queries and keys are encoded in the language space, while values comprise raw sub-trajectories.
@@ 3.1 Background
Decision Transformer. Decision Transformer (Chen et al., 2021, DT) learns a policy from offline data by conditioning on future rewards. This allows rephrasing RL as a sequence modelling problem, where the agent is trained in a supervised manner to map future rewards to actions, often referred to as upside-down RL (Schmidhuber, 2019). To train the DT, we assume access to a pre-collected dataset $\mathcal{D}=\{\tau_{i}\mid 1\leq i\leq N\}$ of $N$ trajectories $\tau_{i}$ that are sampled from the environment via a behavioural policy $\pi_{\beta}$. Each trajectory $\tau\in\mathcal{D}$ consists of state, action, reward, and return-to-go (RTG) quadruplets $\tau_{i}=(s_{0},a_{0},r_{0},\hat{R}_{0},\ldots,s_{T},a_{T},r_{T},\hat{R}_{T})$, where $T$ represents the length of trajectory $\tau_{i}$, and $\hat{R}_{t}=\sum_{t^{\prime}=t}^{T}r_{t^{\prime}}$. The DT $\pi_{\theta}$ is trained to predict the ground truth action $a_{t}$ conditioned on sub-trajectories via cross-entropy or mean-squared error loss, depending on the domain:
@@ 3.2 Retrieval-augmented Decision Transformer (RA-DT)
Figure 2: Illustration of experience reweighting. Given a query trajectory, we retrieve the top $l>k$ most relevant experiences by maximum inner product search. Each experience has an associated task ID and return, based on which we compute its utility. We reweight by $s_{rel}$ and $s_{u}$ to obtain the final retrieval score $s_{ret}$ and return the top-$k$ experiences.
Processing long sequences with DTs is computationally expensive due to the quadratic complexity of the Transformer architecture. To address this challenge, we introduce RA-DT, which equips the DT with an external memory that relies on a vector index for retrieval. Consequently, RA-DT consists of a parametric and a non-parametric component, reminiscent of complementary learning systems (Mcclelland et al., 1995; Kumaran et al., 2016). The former is represented by the DT and learns to predict actions conditioned on the future return. The latter is the retrieval component that searches for relevant experiences, similar to Borgeaud et al. (2022) (see Figure 1).
@@ 3.2.1 Vector Index for Retrieval Augmentation
We aim at augmenting the DT with a vector index (external memory) that allows for the retrieval of relevant experiences. To this end, we build our vector index by leveraging an embedding model $g:\tau\mapsto\mathbb{R}^{d_{r}}$ that takes a trajectory $\tau$ and returns a vector of size $d_{r}$. Given a dataset $\mathcal{D}$ of trajectories, we obtain a set of key-value pairs of our vector index by embedding all sub-trajectories $\tau_{t-C:t}\in\mathcal{D}$ via $g(\cdot)$ to obtain $\mathcal{K}\times\mathcal{V}=\{(g(\tau_{i,t-C:t}),\tau_{i,t-C:t+C})\mid 1\leq i\leq|\mathcal{D}|\}$. Note that values contain sub-trajectories ranging from tCt-C to $t+C$, while keys use sub-trajectories tC:$t$-C:t for a fixed $C$, where $t$ goes over trajectory length in increments of $C$ (see Appendix C.4 for more details). The reason for this choice is that during inference, the model does not have access to future states.
In RAG applications for Natural Language Processing (NLP), a common choice for $g(\cdot)$ is a pre-trained LM. While pre-trained models in NLP are ubiquitous, they are rarely available in RL. A natural choice to instantiate $g(\cdot)$ is to train a DT on the pre-collected dataset $\mathcal{D}$, as they exhibit a well-separated embedding space after pre-training (Schmied et al., 2024). Therefore, they are well suited for retrieval since a new task can be matched to similar tasks in the vector index. As a domain-agnostic alternative, we propose to utilize the FrozenHopfield (FH) mechanism (Paischer et al., 2022) to map trajectories to the embedding space of a pre-trained LM. This enables instantiating $g(\cdot)$ with a pre-trained language encoder. The FH mechanism is parameterized by an embedding matrix $\bm{E}\in\mathbb{R}^{v\times d_{\text{LM}}}$ of a pretrained LM with vocabulary size $v$ and hidden dimension $d_{\text{LM}}$, a random matrix $\bm{P}$ with entries sampled from $\mathcal{N}(0,d_{\text{in}}/d_{\text{LM}})$, and a scaling factor $\beta$ and performs:
@@
We denote $\bm{x}_{t}\in\mathbb{R}^{d_{\text{in}}}$ as the input token and apply the FH position-wise to every state/action/reward token in a sub-trajectory $\tau_{t-C:t}$ separately. Finally, we apply an LM on top of the FH to obtain the keys of our vector index by setting $g(\cdot)=\operatorname{LM}(\operatorname{FH}(\cdot))$. Utilizing the FH enables leveraging the expressive power of frozen pre-trained LMs pre-trained on text as trajectory encoders for RL. This sidesteps the need for pre-training or fine-tuning a domain-specific embedding model and can be incorporated into any existing retrieval-augmentation pipeline.